OpenStudy (anonymous):

solve for A and S for the system a+s=700 10a+7s=5650

OpenStudy (anonymous):

You can use the first equation to build a substitution for a or s, do you know how?

OpenStudy (anonymous):

a+s=700 10a+7s=5650 -10a+-10s=-7000 10a+7s=5650 -3s=-1240 s= some random number that i'm sure is wrong. this is what i've done so far

OpenStudy (anonymous):

You id a good job on everything up to 5650-7000 :) its -1350

OpenStudy (anonymous):

You see, when I did that out the first time I got the -1350 but then this happened -3s=-1350 s=450 a+(450)=700 a= 240 so I added 240 to the 450 to see if it added up to 700, and obviously it added up to 690. I made a mistake in writing it out:/ Thank you so much for helping me see that. Is it possible you could help me with another question? It's very similar to this one.

OpenStudy (anonymous):

Probably, go ahead and post it.

OpenStudy (anonymous):

Just going to warn you, theres a 70% probability I miswrote something and that screwed up my whole answer- this happens alot.

OpenStudy (anonymous):

x+y=900 10x+5y=6850 -5x+-5y=-4500 10x+5y=6850 10x= 2350 x=235 (235)+y=900 y=655 655+235=890

OpenStudy (anonymous):

-5x+-5y=-4500 10x+5y=6850 10x= 2350 check the coefficient on the x again.

OpenStudy (anonymous):

oh. thank you(:

OpenStudy (anonymous):

You've got the right ideas, just be really careful. Your best ally in math is going slow and being careful. You'll always spend less time that way, in the long run. Plus it's far less frustrating. See you!

OpenStudy (anonymous):

Thanks!